# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v1

@@
${}^{1\,}$nstitute for Artificial Intelligence, Peking University
${}^{2\,}$o Lab, Beijing Institute for General Artificial Intelligence
${}^{3\,}$epartment of Automation, Tsinghua University  ${}^{4\,}$hanghai Jiao Tong University
${}^{5\,}$nstitute of Automation, Chinese Academy of Sciences   ${}^{6\,}$niversity of California, Los Angeles
@@
[…]
Figure 1: Comparison of LatentSeek with RL-based fine-tuning and Prompt Engineering. RL-based fine-tuning methods generally require iterative updates to model parameters guided by reward signals. Prompt engineering approaches depend heavily on manually designed prompts. In contrast, LatentSeek performs optimization within the latent space. Of note, the output of LatentSeek may be incoherent and semantically ungrounded (§$\S$3.6).
@@ 1 Introduction
To enhance TTIA performance in reasoning, recent advances Deng et al. (2022), Hao et al. (2024) reveal that reasoning capabilities can be internalized within the latent space [^1] In this work, we take the convention Kong et al. (2025), Hao et al. (2024) that treats the transformers’ output space ahead of the final language model (LM) head as latent space (Figure 1), and the vector in the space as latent representation (Figure 1); refer to §$\S$2.2 for notations. through fine-tuning. However, the training strategies adopted in these works tend to substantially modify the latent space and fail to fully leverage its semantic richness. As a result, their performance remains inferior to that of Chain-of-Thought (CoT). Nevertheless, these studies provide evidence supporting the adequacy of semantic information encoded within latent representations, i.e., the hidden states corresponding to language tokens.
@@
Motivated by these observations, we present the first attempt to perform seeking in the latent space by introducing LatentSeek, a framework that significantly enhances instance-level reasoning at test time. LatentSeek introduces updated instance-specific latent representations that steer the pre-trained model’s reasoning process without modifying its parameters. These latent representations act as a planning or control mechanism that guides the model toward better reasoning paths for each specific problem instance. We optimize latent representations at test time using the policy gradient method Williams (1992) to maximize reward (§$\S$2.3). Specifically, for each reasoning problem, we update the token-wise latent representations using guidance from the reward function, treating them as independent variables. In each iteration, the updated latent representations are decoded into tokens, which serve as inputs for computing the reward. Importantly, the reward function operates in a self-rewarding manner, relying solely on the model’s internal capabilities without incorporating any external information. The process continues until the reward exceeds a predefined threshold or the maximum number of iterations is reached.
@@
Our innovative latent space TTIA method is simple yet surprisingly effective: Notably, LatentSeek achieves an average improvement of 10.75% over the CoT method on the GSM8K dataset, 3.93% on MATH-500, and 4.73% on AIME2024. Furthermore, when using LlaMA3.1-8B-Instruct as the backbone, LatentSeek surpasses prior arts including SimpleRL-Zoo Williams (1992) (+18.1%) and Genius Xu et al. (2025a) (+12.7%), on the GSM8K and MATH-500 datasets. To further assess the potential of the latent space, we conduct idealized experiments utilizing a perfect verifier that relies on ground truth and provides sparse rewards. The experimental results demonstrate a significant performance improvement, with an average gain of 19.12% points over CoT reasoning across all evaluated settings. This highlights the effectiveness of the knowledge elicitation mechanism employed by LatentSeek. Further experiments (see §$\S$3.4) reveal that the model’s performance at test time improves proportionally with the number of update iterations. Notably, when equipped with an ideal verifier, LatentSeek is capable of elevating the performance of a 1.5B parameter model from 54.8% to 82.8% on the MATH-500 dataset, approaching the performance of OpenAI’s o1-preview model. These findings suggest that test-time scaling can be effectively implemented in the latent space, offering a viable alternative to conventional token-space scaling strategiesLiu et al. (2025), Yeo et al. (2025), Xu et al. (2025b).
@@
2 Test-Time Instance-Level Adaptation (TTIA) in Latent Space
2.1 Problem Formulation: TTIA for Reasoning Problems
@@
Given a reasoning problem instance $\mathbf{c}$ as context prompt, a pre-trained auto-regressive language model $\pi$, the probability distribution over a reasoning token sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})$ is:
$\pi(\mathbf{x}|\mathbf{c})=\prod_{t=1}^{T}\pi(x_{t}\mid\mathbf{x}_{<t},\mathbf{c}),$ (1)
@@
where $\mathbf{x}_{<t}$ denotes the sequence of tokens preceding position $t$. The objective of solving for $\mathbf{c}$ is to find a reasoning sequence $\mathbf{x}$, such that the final answer $a\sim\pi(\cdot|\mathbf{x},\mathbf{c})$, generated following Equation 1, is consistent with the ground truth. However, at test time, the ground truth is unknown, and thus a reward function $R(\mathbf{x},\mathbf{c})$ is introduced to evaluate the reasoning token sequence. Concluding the above, the TTIA objective for a reasoning problem is as follows:
@@
Please refer to Appendix B for examples.
2.2 TTIA with Policy Gradient in Latent Space
To solve the problem in Equation 2, we reformulate the task as optimizing over a sequence of latent representations rather than directly searching for reasoning tokens. Specifically, for a given input sequence $\mathbf{x}$, we denote a corresponding sequence of latent representations $\mathbf{z}=(z_{1},z_{2},\ldots,z_{T})$, where $z_{t}\in\mathbb{R}^{d}$ lies in the latent space of $x_{t}$. To identify the optimal sequence of latent representations, TTIA in the latent space aims to optimize the following objective:
$\mathbf{z}^{*}=\arg\max_{\mathbf{z}}\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}|\mathbf{z})}[R(\mathbf{x},\mathbf{c})].$ (3)
Based on this formula, we will present the sampling strategy and optimization procedure.
Independent Sampling.
For concrete calculation of $\mathbf{x}\sim\pi(\mathbf{x}|\mathbf{z})$, we start by pointing out that the tokens are conditionally independent given their respective latent representations:
$\pi(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}\pi(x_{t}|z_{t}),$ (4)
which implies that the sampling can be carried out on a token-by-token basis.
@@ 2.2 TTIA with Policy Gradient in Latent Space
We optimize the latent representations using a direct policy gradient approach based on REINFORCE (Williams, 1992). Assuming the independence of the latent representations (see Appendix A for a theoretical justification), the update process is:
@@
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}|\mathbf{z})}\left[R(\mathbf{x},\mathbf{c})\nabla_{\mathbf{z}}\log\pi(\mathbf{x}|\mathbf{z})\right].$ (6)
Leveraging Equation 4, the gradient of the $t$-th latent representation is:
$[\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})]_{t}=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}|\mathbf{z})}\left[R(\mathbf{x},\mathbf{c})\nabla_{z_{t}}\log\pi(x_{t}|z_{t})\right],$ (7)
where the expectation is approximated using the empirical mean in practical implementation.
@@ 2.3 LatentSeek Algorithm
Problem $\mathbf{c}$, learning rate $\eta$, pre-trained model $\pi$, reward threshold $\tau$, fraction ratio $\rho\in(0,1]$
$\mathbf{z}^{(0)}\leftarrow\pi(\mathbf{c})$triangleright Initial latent representations before the LM head.
$r,\mathbf{x}\leftarrow R(\mathbf{x},\mathbf{c})$triangleright Self Reward with Equation 10 and the final answer.
@@
$\mathbf{z}^{(0)}\leftarrow[z_{1}^{(0)},z_{2}^{(0)},\ldots,z_{\rho T}^{(0)}]$triangleright Keep a $\rho$ fraction, such as 20%
while $k\leq K$ and $r\leq\tau$ do $\triangleright$ $K$ is typically $10$
     $\mathbf{z}^{(k)}\leftarrow\mathbf{z}^{(k-1)}+\eta\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z}^{(k-1)})$triangleright Independently update. $\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z}^{(k-1)})$: Equation 7.
     $x_{t}\sim\pi(x_{t}|z^{(k)}_{t}),\,t=1,2,\ldots,\rho\cdot T$triangleright Decode tokens: Equation 8.
     $r,\tilde{\mathbf{x}}\leftarrow R([x_{1},x_{2},\ldots,x_{\rho T}],\mathbf{c})$triangleright Self Reward with Equation 10 and the final answer.
@@
The LatentSeek algorithm is described in Algorithm 1. This algorithm iteratively refines the latent representations based on the rewards of generated reasoning paths, effectively performing a guided search through the reasoning space specific to the given problem instance. After each refinement step, the latent representations are decoded into tokens to calculate a reward signal. This signal is then employed to direct the search process in the subsequent iteration. Along with the reward signal, the final output $\tilde{\mathbf{x}}$ is also explicitly provided. The process runs for a small number of iterations (typically 2-10), stopping early if the reward exceeds a threshold.
Decoding Strategy.
Greedy decoding is employed as an approximation method for sampling to enhance computational efficiency:
$x_{t}=\arg\max_{v\in V}\pi(v|z_{t}),$ (8)
where $V$ denotes the vocabulary set, $t$ represents the token position.
Reward Function.
The reward function $R(\mathbf{x},\mathbf{c})$ is defined as follows: Given a sequence of decoded tokens $\mathbf{x}$, the pre-trained model is guided to generate a complete sequence $\tilde{\mathbf{x}}$, which encapsulates the final answer. This sequence is obtained by greedy decoding of autoregressive generation:
$\tilde{\mathbf{x}}=\arg\max_{\tilde{\mathbf{x}}}\pi(\tilde{\mathbf{x}}\mid\mathbf{x},\mathbf{c}).$ (9)
Subsequently, the sequence $\tilde{\mathbf{x}}$ is evaluated using a self-reward prompt to compute the reward:
$R(\mathbf{x},\mathbf{c})\sim\pi(\cdot\mid\tilde{\mathbf{x}},\mathbf{c},\text{prompt}_{\text{self-reward}}).$ (10)
Noticed that, in the algorithm, the complete answer $\tilde{\mathbf{x}}$ is also returned.
@@
To ensure more stable and efficient optimization, we propose the integration of two complementary techniques. First, CoT Initialization: The initial latent representation, denoted as $\mathbf{z}^{(0)}$, is derived from the CoT reasoning sequence. This approach leverages the reasoning capabilities of CoT to establish an effective starting point for subsequent optimization. Second, Fractional Sequence Optimization: Instead of optimizing the entire sequence of latent representations $\mathbf{z}=[z_{1},z_{2},\ldots,z_{T}]$, we propose to optimize only a subsequence $[z_{1},z_{2},\ldots,z_{\rho T}]$, where $\rho\in(0,1]$ is a hyperparameter. This strategy reduces computational cost and promotes stable optimization through the careful selection of $\rho$. The hyperparameter $\rho$ must strike a balance between two competing objectives: maintaining adequate representational capacity to support effective exploration, and limiting the number of latent representations being updated. Excessive modification of latent representations—especially when their decoded outputs lack semantic coherence—can compromise the reliability of the reward function.
@@ 3 Empirical Results
In this section, we present our experimental setup, results, and analysis. Additional implementation details are provided in Sections C.1, C.3, C.2, C.4, C.5 and C.6.
@@ 3.1 Experimental Setup
We employ the mathematical reasoning prompts introduced by Lifshitz et al. (2025) to elicit the model’s self-reward computation. The same prompt structure is applied consistently across all tasks and models. For specific tasks and models, we further incorporate a format-based reward DeepSeek-AI (2025). To further illustrate the potential of our paradigm, we introduce a Perfect Sparse Reward Model (PSRM), the details of which are discussed in Section 3.3.
@@
For robustness consideration, we use two prompts for evaluation, with the first requiring a wrapped answer with \$boxed\{\}$ Yang et al. (2024), Team (2024a) (prompt 1) and the second asking to format the answer as a json (prompt 2).

### v2

@@
$^{1\,}$Institute for Artificial Intelligence, Peking University
$^{2\,}$NLCo Lab, Beijing Institute for General Artificial Intelligence
$^{3\,}$Department of Automation, Tsinghua University  $^{4\,}$Shanghai Jiao Tong University
$^{5\,}$Institute of Automation, Chinese Academy of Sciences   $^{6\,}$University of California, Los Angeles
@@
Reasoning ability, a core component of human intelligence, continues to pose a significant challenge for Large Language Models (LLMs) in the pursuit of AGI. Although model performance has improved under the training scaling law, significant challenges remain, particularly with respect to training algorithms—such as catastrophic forgetting—and the limited availability of novel training data. As an alternative, test-time scaling enhances reasoning performance by increasing test-time computation without parameter updating. Unlike prior methods in this paradigm focused on token space, we propose leveraging latent space for more effective reasoning and better adherence to the test-time scaling law. We introduce LatentSeek, a novel framework that enhances LLM reasoning through Test-Time Instance-level Adaptation (TTIA) within the model’s latent space. Specifically, LatentSeek leverages policy gradient to iteratively update latent representations, guided by self-generated reward signals. LatentSeek is evaluated on a range of reasoning benchmarks, including GSM8K, MATH-500, and AIME2024, across multiple LLM architectures. Results show that LatentSeek consistently outperforms strong baselines, such as Chain-of-Thought prompting and fine-tuning-based methods. Furthermore, our analysis demonstrates that LatentSeek is highly efficient, typically converging within a few iterations for problems of average complexity, while also benefiting from additional iterations, thereby highlighting the potential of test-time scaling in the latent space. These findings position LatentSeek as a lightweight, scalable, and effective solution for enhancing the reasoning capabilities of LLMs. Code https://github.com/bigai-nlco/LatentSeek Project https://bigai-nlco.github.io/LatentSeek/
Figure 1: Comparison of LatentSeek with RL-based fine-tuning and Prompt Engineering. RL-based fine-tuning methods generally require iterative updates to model parameters guided by reward signals. Prompt engineering approaches depend heavily on manually designed prompts. In contrast, LatentSeek performs optimization within the latent space. Of note, the output of LatentSeek may be incoherent and semantically ungrounded ($\S$3.6).
@@ 1 Introduction
A prevalent approach in TTIA is prompt engineering, which suffers from limited expressiveness. To enhance TTIA performance in reasoning, recent advances Deng et al. (2022), Hao et al. (2024) reveal that reasoning capabilities can be internalized within the latent space [^1] In this work, we take the convention Kong et al. (2025), Hao et al. (2024) that treats the transformers’ output space ahead of the final language model (LM) head as latent space (Figure 1), and the vector in the space as latent representation (Figure 1); refer to $\S$2.2 for notations. through fine-tuning. However, the training strategies adopted in these works tend to substantially modify the latent space and fail to fully leverage its semantic richness. As a result, their performance remains inferior to that of Chain-of-Thought (CoT). Nevertheless, these studies provide evidence supporting the adequacy of semantic information encoded within latent representations, i.e., the hidden states corresponding to language tokens.
@@
Motivated by these observations, we present the first attempt to perform seeking in the latent space by introducing LatentSeek, a framework that significantly enhances instance-level reasoning at test time. LatentSeek introduces updated instance-specific latent representations that steer the pre-trained model’s reasoning process without modifying its parameters. These latent representations act as a planning or control mechanism that guides the model toward better reasoning paths for each specific problem instance. We optimize latent representations at test time using the policy gradient method Williams (1992) to maximize reward ($\S$2.3). Specifically, for each reasoning problem, we update the token-wise latent representations using guidance from the reward function, treating them as independent variables. In each iteration, the updated latent representations are decoded into tokens, which serve as inputs for computing the reward. Importantly, the reward function operates in a self-rewarding manner, relying solely on the model’s internal capabilities without incorporating any external information. The process continues until the reward exceeds a predefined threshold or the maximum number of iterations is reached.
@@
Our innovative latent space TTIA method is simple yet surprisingly effective: Notably, LatentSeek achieves an average improvement of 10.75% over the CoT method on the GSM8K dataset, 3.93% on MATH-500, and 4.73% on AIME2024. Furthermore, when using LlaMA3.1-8B-Instruct as the backbone, LatentSeek surpasses prior arts including SimpleRL-Zoo Williams (1992) (+18.1%) and Genius Xu et al. (2025a) (+12.7%), on the GSM8K and MATH-500 datasets. To further assess the potential of the latent space, we conduct idealized experiments utilizing a perfect verifier that relies on ground truth and provides sparse rewards. The experimental results demonstrate a significant performance improvement, with an average gain of 19.12% points over CoT reasoning across all evaluated settings. This highlights the effectiveness of the knowledge elicitation mechanism employed by LatentSeek. Further experiments (see $\S$3.4) reveal that the model’s performance at test time improves proportionally with the number of update iterations. Notably, when equipped with an ideal verifier, LatentSeek is capable of elevating the performance of a 1.5B parameter model from 54.8% to 82.8% on the MATH-500 dataset, approaching the performance of OpenAI’s o1-preview model. These findings suggest that test-time scaling can be effectively implemented in the latent space, offering a viable alternative to conventional token-space scaling strategiesLiu et al. (2025), Yeo et al. (2025), Xu et al. (2025b).
@@
2 Test-Time Instance-Level Policy Gradient in Latent Space
2.1 Problem Formulation: Test-Time Instance-Level Reasoning
@@
Let $\mathbf{c}$ be the context prompt of a reasoning problem instance and $\pi$ a pre-trained auto-regressive language model (LM) composed of a Transformer backbone $\pi_{\text{Transformer}}$ (Vaswani et al., 2017) and a LM head $\pi_{\text{LM-head}}$. For a reasoning sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})$,
$\pi(\mathbf{x}\mid\mathbf{c})=\prod_{t=1}^{T}\pi(x_{t}\mid\mathbf{x}_{<t},\mathbf{c}),\quad\quad\pi(x_{t}\mid\mathbf{x}_{<t},\mathbf{c})=\pi_{\text{LM-head}}(x_{t}\mid z_{t}),$ (1)
@@
where $\mathbf{x}_{<t}$ denotes the sequence of tokens preceding position $t$ and $z_{t}:=\pi_{\text{Transformer}}(\mathbf{x}_{<t},\mathbf{c})$ is the latent representation associated with $x_{t}$. Following Hao et al. (2024), we refer to $z_{t}$ as lying in the latent space of $x_{t}$. At test time, the ground truth is unknown, and thus a reward function $R(\mathbf{x},\mathbf{c})$ is introduced to evaluate the reasoning token sequence. Concluding the above, the objective for a reasoning problem is to find an optimal reasoning path:
@@
Please refer to Appendix˜B for examples.
2.2 Reasoning via Policy Gradient in Latent Space
To solve the problem in Equation˜2, we reformulate the task as optimizing over a sequence of latent representations rather than directly searching for tokens. Specifically, we denote a sequence of latent representations $\mathbf{z}=(z_{1},z_{2},\ldots,z_{N})$, where $z_{t}$ lies in the latent space of $x_{t}$ and $N$ is typically an integer smaller than or equal to $T$. To identify the optimal sequence of latent representations, the to-be-optimized objective is as follows:
$\mathbf{z}^{*}=\operatorname*{arg\,max}_{\mathbf{z}}\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}[R(\mathbf{x},\mathbf{c})].$ (3)
To sample $\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})$, we first decode the latents $\mathbf{z}$ into their corresponding tokens, then continue with autoregressive generation, since the complete sequence may extend beyond the latent-initialized reasoning path. As $T$ is the complete sequence length, the factorization is
$\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})=\underbrace{\prod_{t=1}^{N}\pi_{\text{LM-head}}(x_{t}\mid z_{t})}_{\text{decode from latents}}\underbrace{\prod\limits_{t=N+1}^{T}\pi(x_{t}\mid\mathbf{x}_{<t},\mathbf{c})}_{\text{continue generation}},$ (4)
where, if the generation stops at $t=N$, the second product is empty and equals 1.
@@ 2.2 Reasoning via Policy Gradient in Latent Space
Given the objective in Equation˜3, we optimize the latent representations using a direct policy gradient approach based on REINFORCE (Williams, 1992). Assuming the independence of the latent representations, the update process is:
@@
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}\left[R(\mathbf{x},\mathbf{c})\nabla_{\mathbf{z}}\log\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})\right].$ (6)
Leveraging Equation˜4, the gradient of the $t$-th latent representation is:
$\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})]_{t}=\mathbb{E}_{\mathbf{x}\sim\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})}\left[R(\mathbf{x},\mathbf{c})\nabla_{z_{t}}\log\pi(x_{t}\mid z_{t})\right],$ (7)
where the expectation is approximated using the empirical mean in practical implementation. Please refer to Appendices˜C and D for a theoretical justification of this assumption and a detailed derivation of the policy-gradient update.
@@ 2.3 LatentSeek Algorithm
Problem $\mathbf{c}$, learning rate $\eta$, pre-trained model $\pi$, reward threshold $\tau$, fraction ratio $\rho$, max iterations $K$
$\mathbf{x},\mathbf{z}\leftarrow\pi(\mathbf{x}\mid\mathbf{c})$ $\triangleright$ Initialize latent representations with CoT: Equation˜1.
$r\leftarrow R(\mathbf{x},\mathbf{c})$ $\triangleright$ Self Reward with Equation˜8.
@@
$\mathbf{z}\leftarrow[z_{1},z_{2},\ldots,z_{\rho T}]$ $\triangleright$ Keep a $\rho$ fraction, such as 20%
while $k\leq K$ and $r\leq\tau$ do $\triangleright$ $K$ is typically 1010
  $\mathbf{z}\leftarrow\mathbf{z}+\eta\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})$ $\triangleright$ Update $\nabla_{\mathbf{z}}\mathcal{J}(\mathbf{z})$: Equations˜6 and 7.
  $\mathbf{x}\sim\pi(\mathbf{x}|\mathbf{z},\mathbf{c})$ $\triangleright$ Sample sequence: Equation˜4.
  $r\leftarrow R(\mathbf{x},\mathbf{c})$ $\triangleright$ Self Reward with Equation˜8.
@@
The LatentSeek algorithm is summarized in Algorithm˜1. At a high level, it performs an instance-level, gradient-guided search over a latent space by iteratively refining latent representations and querying a self-rewarding mechanism. Concretely, at iteration $k$, the current latents $\mathbf{z}$ are decoded to a full sequence $\mathbf{x}$ through a sampling method. Conditioned on $\mathbf{x}$ and the instance context $\mathbf{c}$, a scalar reward is then obtained from a self-rewarding mechanism:
$R(\mathbf{x},\mathbf{c})\sim\pi(\cdot\mid\mathbf{x},\mathbf{c},\text{prompt}_{\text{self-reward}}).$ (8)
This reward signal drives a gradient-guided refinement step that updates the latent representations, thereby biasing subsequent decodes toward higher-reward reasoning paths. The process runs for a small number of iterations (typically less than 3), stopping early if the reward exceeds a threshold.
@@
To ensure more stable and efficient optimization, we propose the integration of two complementary techniques. First, CoT Initialization: The initial latent representation is derived from the CoT reasoning sequence. This approach leverages the reasoning capabilities of CoT to establish an effective starting point for subsequent optimization. Second, Fractional Sequence Optimization: Instead of optimizing the entire sequence of latent representations $\mathbf{z}=[z_{1},z_{2},\ldots,z_{T}]$, we propose to optimize only a subsequence $z_{1},z_{2},\ldots,z_{\rho T}]$, where $\rho\in(0,1]$ is a hyperparameter. The hyperparameter $\rho$ must strike a balance between two competing objectives: maintaining adequate representational capacity to support effective exploration, and limiting the number of latent representations being updated. Excessive modification of latent representations—especially when their decoded outputs lack semantic coherence—can compromise the reliability of the reward function.
@@ 3 Empirical Results
In this section, we present our experimental setup, results, and analysis. Additional implementation details are provided in Sections˜F.1, F.3, F.2, F.4, F.5 and F.6.
@@ 3.1 Experimental Setup
We employ the mathematical reasoning prompts introduced by Lifshitz et al. (2025) to elicit the model’s self-reward computation. The same prompt structure is applied consistently across all tasks and models. For specific tasks and models, we further incorporate a format-based reward DeepSeek-AI (2025). To further illustrate the potential of our paradigm, we introduce a Perfect Sparse Reward Model (PSRM), the details of which are discussed in Section˜3.3.
@@
[…]